# Jet schemes, Newton polygons and continued fractions

### v2

@@ 7. Main Theorems
In this section, we describe the irreducible components of the jet schemes over the origin of a Newton non-degenerate plane curve singularity. First, let’s present some definitions and notations.
Let $f\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. A tropical decomposition of $f$ is $f=f_{1}...f_{t}$ where $T(f_{i})$ is a semi-line in $\mathbf{R}_{\geq 0}^{2}$ passing through the origin and $T(f_{i})\cap T(f_{j})=\{(0,0)\}$ for $1\leq i<j\leq t$.
Denote now by $\sigma=Conv[T(f)]$. Introduce the notations
@@
where $B^{*}_{m}$ is obtained by removing from $B_{m}$ at $m=\nu_{\alpha}(f)$ every $\alpha+(1,1)$ : $\alpha\in B_{m-1}\cap\partial\sigma$. Theorem 7 gives us the decomposition of $\mathcal{C}^{0}_{m}$ into irreducible components.
@@
Let $f=f_{1}...f_{t}\in\mathcal{O}_{k^{2}}$ be a tropical decomposition of a Newton non-degenerate polynomial. Decompose $f_{i}=\prod_{j=1}^{r_{i}}f_{(i,j)}$ into irreducible factors in $\mathcal{O}_{k^{2}}$. Then, we have the decomposition into irreducible components
@@
Recall for $v\in\{x,y\}$: $Cont^{>0}_{\mathcal{C}}(v)_{m}=\bigsqcup_{i=1}^{m}Cont^{i}_{\mathcal{C}}(v)_{m}$ where we set
$Cont^{i}_{\mathcal{C}}(v)_{m}=\{v^{(i)}\neq 0\}\cap\{v^{(0)}=...=v^{(i-1)}=0\}\cap\{f^{(0)}=...=f^{(m)}=0\}$
Set also, $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=Cont^{i}_{\mathcal{C}}(x)_{m}\cap Cont^{j}_{\mathcal{C}}(y)_{m}$. Start with the observation,
$\mathcal{C}^{0}_{m}=\mathcal{H}^{(1,1)}_{m}\cap F_{m}=Cont^{>0}_{\mathcal{C}}(x)_{m}\cap Cont_{\mathcal{C}}^{>0}(y)_{m}=\bigsqcup_{1\leq i,j\leq m}Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}$
Then, if we set $\alpha=(i,j)$ and notice that $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=F^{(i,j)}_{m}\cap\mathbf{A}^{2}_{m}$, thus,
@@
$A_{m}=\bigcup_{i=1}^{t}A_{m}(i)$
If $\alpha\notin A_{m}$ then $f^{[\nu_{\alpha}(f)]}=c\cdot{x^{(\alpha_{1})}}^{k_{1}}{y^{(\alpha_{2})}}^{k_{2}}\mod\alpha$ for some constant $c$. As $x^{(\alpha_{1})}$ and $y^{(\alpha_{2})}$ are $\neq 0$ in $F^{\alpha}_{m}$ then $F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m}=\varnothing$ if $\alpha\notin A_{m}$. Now, we can write:
$\mathcal{C}^{0}_{m}=\bigsqcup_{\alpha\in\Omega^{>}}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})\sqcup\bigsqcup_{i=1}^{t}\bigsqcup_{\alpha\in A_{m}(i)}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})$
Now if $\alpha\in A_{m}(i)$ then $F^{\alpha}_{m}=F^{\alpha}_{i,k}$ where $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$,
$F^{\alpha}_{i,k}=\bigsqcup_{j=1}^{r_{i}}F^{\alpha}_{(i,j),k_{ij}}$
because of lemma 2. Now, as $\mathcal{H}^{\alpha}_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for $\nu_{\alpha}(f)>m$ then we obtain after ordering along $\leq\times\leq$ and using theorem 4,
The irreducibility of these components follow from 5, 6 and 4. ∎
We have just proven that $Jet(f)$ has two kinds of vertices: vertices of the form $\mathcal{H}^{\gamma}_{m}$ and others that do not vanish and continue to $\infty$. The below theorem characterizes them.
@@
In this proof, set $c=codim(V_{m};\mathbf{A}^{2}_{m})$. If $V_{m}=\mathcal{H}^{\gamma}_{m}$ then $c=|\gamma|$. Now, let’s show $e+c=2(m+1)$.
$\mathcal{O}_{V_{m}}=\mathcal{O}_{\mathbf{A}^{2}_{m}}/\langle x^{(0)},...,x^{(\gamma_{1}-1)},y^{(0)},...,y^{(\gamma_{2}-1)}\rangle$
$e\in\mathbf{N}$ is the least integer such that we have a surjective map of rings
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}$
Thus, $e=dim(\mathbf{A}^{2}_{m})-|\gamma|=2(m+1)-|\gamma|$. Thus, $e+c=2(m+1)$ ie $e=d$. Now, if $V_{m}$ isn’t of the form $\mathcal{H}^{\gamma}_{m}$ then there exists $\alpha\in T(f_{ij})$ for some $i,j$ such that
@@
with $m\geq\nu_{\alpha}(f)$. Also, $c=|\alpha|+1+m-\nu_{\alpha}(f)$, there are surjective ring maps:
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}\rightarrow\mathcal{O}_{\mathcal{H}^{\alpha}_{m}}$
Thus, $e\geq 2(m+1)-|\alpha|$. Then, we obtain:
$e+c\geq 2(m+1)-|\alpha|+c=2(m+1)+1+m-\nu_{\alpha}(f)$
Thus, $e+c>2(m+1)$. Thus, $e>d$. This completes the proof.
Now, we show that $Jet(f)$ determines the embedded topological type of the singularity of the Newton non-degenerate plane curve defined by $f$.
@@
Let $f,g\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. If the jet-components graph $Jet(f)=Jet(g)$ then $f$ and $g$ have the same embedded topological type.
@@
According to theorem 8, $J:=Jet(f)$ determines $B^{*}:=\bigcup_{m\in\mathbf{N}_{0}}B^{*}_{m}$. In other words, $B^{*}_{m}:=B^{*}_{m}(f)=B^{*}_{m}(g)$ for all $m$. Denote by $R^{*}$ the induced subgraph of $J$ obtained by removing the vertices of the form $\mathcal{H}^{\gamma}_{m}$ : $\gamma\in B^{*}_{m}$, $\nu_{\gamma}(f)>m\in\mathbf{N}_{0}$ and $\mathcal{H}^{\gamma}_{m}$ could not be joined with a vertex outside of $B^{*}$. Define $A=\{\alpha:\exists m:\mathcal{H}^{\alpha}_{m}\in R^{*}\}$. Note that:
$\alpha\in A\Leftrightarrow\mathcal{H}^{\alpha}_{\nu_{\alpha}(f)-1}\in R^{*},\alpha\in T(f)\cap\mathbf{N}^{2}$
Set $T=\bigcup_{\alpha\in A}\mathbf{R}_{\geq 0}\cdot\alpha$ which is then the tropical variety of both $f$ and $g$. Now, let’s denote $f=f_{1}...f_{r}$ and $g=g_{1}...g_{s}$ as product of branches. Let $\alpha_{i}\in A$. Denote by $T_{\alpha_{i}}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ the tropical variety of the branches $f_{1},...,f_{r_{i}},g_{1},...,g_{s_{i}}$ with $r_{i}\leq r$ and $s_{i}\leq s$. Since $f$ and $g$ are Newton non-degenerate with the same graph $J$ then the there are exactly $r_{i}=s_{i}$ vertices above $\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(f)-1}=\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(g)-1}$ connected to it. Thus, $r_{i}=s_{i}$ and $r=s$. Also, for $1\leq i<j\leq l<\ell\leq r$,
$(f_{i},f_{j})_{0}=\nu_{\alpha_{i}}(f_{i})=\nu_{\alpha_{i}}(f_{j})$
$(f_{i},f_{\ell})_{0}=\nu_{\alpha_{i}}(f_{\ell})$
The same for $g$. Denote by $F=f/(f_{1}...f_{l})$ and $G=g/(g_{1}...g_{l})$. So,
$l\nu_{\alpha_{i}}(f_{1})+\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)=l\nu_{\alpha_{i}}(g_{1})+\nu_{\alpha_{i}}(G)$
$\nu_{\alpha_{i}}(f_{1})=lcm(\alpha_{i,1},\alpha_{i,2})=\nu_{\alpha_{i}}(g_{1})$. Since $\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)$ then $\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(G)$. This holds for all $\alpha_{i}\in A$. So, we have shown that $f$ and $g$ have the same number of branches, the same intersection multiplicities (of their branches) and the same tropical varieties for branches (since $r_{i}=s_{i}$). Now using theorem 3, $f$ and $g$ have the same embedded topological type. ∎
As we did in Figures 1,3 and 4, we can associate to $J=Jet(f)$ a special staircase walk $J_{SC}(f)$ in $\mathbf{R}^{2}_{\geq 0}$. Retake the same notations of theorem 9. Denote by $T_{i}=T(f_{i})$ where we suppose $T_{j}$ above $T_{i}$ for $i<j$. Let’s define a dictionnary for this walk: every drawn $\alpha\in\mathbf{N}^{2}$, represents the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$ of $\mathcal{C}^{0}_{m}$ : $m<\nu_{\alpha}(f)$. The red arrows correspond to irreducible components of $\mathcal{C}^{0}_{m}$ that don’t vanish and continue till $\infty$.
@@
    If $t=1$, then the staircase walk is identical to that of Figure 3 except that at each $\alpha\in T(f)$ we add $r$ red arrows above $\alpha$.
    Here is a more detailed description of it: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$ and recall the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from the theorem 1:
    • Start at $(1,1)$ ie $\lambda=(0,0)$ here;
    • Move $d_{1}$ vertical unit steps followed by one horizontal step;
    • Move $d_{2}$ vertical unit steps followed by one horizontal step;
    • Move $d_{p-1}$ vertical unit steps followed by one horizontal step;
    • Finally, move $d_{p}-1$ vertical steps; we reach $(p,q)$;
    • Add this time $r$ red arrow above $(p,q)$;
    • Take $\lambda=(p,q)$ and start again;
    Repeat the above steps whenever we reach $\lambda\in L_{(p,q)}\cap\mathbf{N}_{0}^{2}$ till reaching $\lambda+(p,q)$ above which we add $r$ red arrows.
    Note that each time we move from a point $\alpha$ to $\alpha+(0,1)$ or $\alpha+(1,0)$, we delete $\alpha$ and we replace it by the new point. The red arrows are never deleted.
@@
    If $t=2$, then set $\sigma=T_{1}+T_{2}$; the staircase walk obeys to the rules:
    • Start always at $(1,1)$;
    • If $\alpha\in\overset{\circ}{\sigma}$, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
    • If $\alpha\in T_{1}$, add $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0)$;
    • If $\alpha$ is above $T_{2}$ replace $\alpha$ by $\alpha+(1,0)$;
    • If $\alpha\in T_{2}$, add $r_{2}$ red arrows above $\alpha$ and add $\alpha+(0,1)$;
    • If $\alpha$ is below $T_{1}$ replace $\alpha$ by $\alpha+(0,1)$.
@@
    If $t>2$, set $T=T(f)$, $\sigma=T_{1}+T_{t}$; the staircase walk obeys to the rules:
    • Start always at $(1,1)$;
    • If $\T\alpha\in\sigma\backslash$ T, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
    • If $\alpha\in T_{i}$ : $2\leq i\leq t-1$ add $r_{i}$ red arrows above $\alpha$ and add $\alpha+(1,0)$ with $\alpha+(0,1)$;
    • If $\alpha\in T_{1}$, add $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0);$
    • If $\alpha$ is above $T_{t}$ replace $\alpha$ by $\alpha+(1,0)$;
    • If $\alpha\in T_{t}$, add $r_{t}$ red arrows above $\alpha$ and add $\alpha+(0,1)$;
    • If $\alpha$ is below $T_{1}$ replace $\alpha$ by $\alpha+(0,1)$.
@@
$\mathbf{A}^{2}_{m}\cap F^{\alpha}_{ij,k}:\alpha\in T(f_{i})$ Nonvanishing $(\alpha,ij,\infty)$
Table 3. Dictionary of notations for $f$ Newton non-degenerate.
@@
For $\mathcal{C}=\{f=0\}\subset k^{2}$ a Newton plane curve, we can extract the irreducible components of $\mathcal{C}^{0}_{m}$ for all $m\in\mathbf{N}$ from $J_{SC}(f)$.
@@
Consider first an edge and denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$. Now consider a nonvanishing component represented by $(\alpha,ij,\infty)$. Add on the arrow above $\alpha$ a countably infinite amount of vertices. These represent the nonvanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$. The obtained graph is the jet components graph. ∎

### v3

@@ 7. Main Theorems
In this section, we describe the irreducible components of the jet schemes over the origin of a Newton non-degenerate plane curve singularity. First, we fix some notation and present some basic definitions.
Let $f\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. A tropical decomposition of $f$ is $f=f_{1}...f_{t}$ where $T(f_{i})$ is a semi-line in $\mathbf{R}_{\geq 0}^{2}$ passing through the origin and $T(f_{i})\cap T(f_{j})=\{(0,0)\}$ for $1\leq i<j\leq t$. Next, let $\sigma=Conv[T(f)]$, and define
@@
where $B^{*}_{m}$ is obtained from $B_{m}$ by removing at $m=\nu_{\alpha}(f)$, every $\alpha+(1,1)$ with $\alpha\in B_{m-1}\cap\partial\sigma$.
The next theorem gives us the decomposition of $\mathcal{C}^{0}_{m}$ into irreducible components.
@@
Let $f=f_{1}...f_{t}\in\mathcal{O}_{k^{2}}$ be a tropical decomposition of a Newton non-degenerate polynomial. Decompose $f_{i}=\prod_{j=1}^{r_{i}}f_{(i,j)}$ into irreducible factors in $\mathcal{O}_{k^{2}}$. Then, we have the following decomposition of $\mathcal{C}^{0}_{m}$ into irreducible components
@@
Recall that for $v\in\{x,y\}$, we know that $Cont^{>0}_{\mathcal{C}}(v)_{m}=\bigsqcup_{i=1}^{m}Cont^{i}_{\mathcal{C}}(v)_{m}$ where
$Cont^{i}_{\mathcal{C}}(v)_{m}=\{v^{(i)}\neq 0\}\cap\{v^{(0)}=...=v^{(i-1)}=0\}\cap\{f^{(0)}=...=f^{(m)}=0\}.$
Now setting $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=Cont^{i}_{\mathcal{C}}(x)_{m}\cap Cont^{j}_{\mathcal{C}}(y)_{m}$, and observing that
$\mathcal{C}^{0}_{m}=\mathcal{H}^{(1,1)}_{m}\cap F_{m}=Cont^{>0}_{\mathcal{C}}(x)_{m}\cap Cont_{\mathcal{C}}^{>0}(y)_{m}=\bigsqcup_{1\leq i,j\leq m}Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m},$
and $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=F^{(i,j)}_{m}\cap\mathbf{A}^{2}_{m}$ for $\alpha=(i,j)$, then clearly we obtain
@@
$A_{m}=\bigcup_{i=1}^{t}A_{m}(i).$
If $\alpha\notin A_{m}$ then $f^{[\nu_{\alpha}(f)]}=c\cdot{x^{(\alpha_{1})}}^{k_{1}}{y^{(\alpha_{2})}}^{k_{2}}\mod\alpha$ for some constant $c$. As $x^{(\alpha_{1})}$ and $y^{(\alpha_{2})}$ are $\neq 0$ in $F^{\alpha}_{m}$ then $F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m}=\varnothing$ for $\alpha\notin A_{m}$. Note that we can write:
$\mathcal{C}^{0}_{m}=\bigsqcup_{\alpha\in\Omega^{>}}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})\sqcup\bigsqcup_{i=1}^{t}\bigsqcup_{\alpha\in A_{m}(i)}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m}).$
In view of Lemma 2, we know that if $\alpha\in A_{m}(i)$ then $F^{\alpha}_{m}=F^{\alpha}_{i,k}$ where $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$, and
$F^{\alpha}_{i,k}=\bigsqcup_{j=1}^{r_{i}}F^{\alpha}_{(i,j),k_{ij}}.$
Now, as $\mathcal{H}^{\alpha}_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for $\nu_{\alpha}(f)>m$, then after ordering along $\leq\times\leq$ and using Theorem 4, we obtain:
$\mathcal{C}^{0}_{m}=\bigcup_{\gamma\in B^{*}_{m}}\mathcal{H}^{\gamma}_{m}\cup\bigcup_{i=1}^{t}\bigcup_{\alpha\in A_{m}(i)}\bigcup_{j=1}^{r_{i}}\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{(i,j),k_{ij}}}.$
The irreducibility of these components follow then from Lemma 5, Lemma 6 and Theorem 5. ∎
It is worth mentioning here that we have just shown that $Jet(f)$ has two kinds of vertices: vertices of the form $\mathcal{H}^{\gamma}_{m}$ and others that do not vanish and continue to $\infty$. The next theorem deals with the characterization of such vertices.
@@
First, we set $c=codim(V_{m};\mathbf{A}^{2}_{m})$. If $V_{m}=\mathcal{H}^{\gamma}_{m}$ then $c=|\gamma|$. Next, we show that $e+c=2(m+1)$. For,
$\mathcal{O}_{V_{m}}=\mathcal{O}_{\mathbf{A}^{2}_{m}}/\langle x^{(0)},...,x^{(\gamma_{1}-1)},y^{(0)},...,y^{(\gamma_{2}-1)}\rangle,$
and $e\in\mathbf{N}$ is the least integer such that we have a surjective map of rings
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}.$
Hence, $e=dim(\mathbf{A}^{2}_{m})-|\gamma|=2(m+1)-|\gamma|$. Thus, $e+c=2(m+1)$ i.e. $e=d$. Now, if $V_{m}$ is not of the form $\mathcal{H}^{\gamma}_{m}$ then there exists $\alpha\in T(f_{ij})$ for some $i,j$ such that
@@
with $m\geq\nu_{\alpha}(f)$. Also, $c=|\alpha|+1+m-\nu_{\alpha}(f)$, then there are surjective ring maps:
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}\rightarrow\mathcal{O}_{\mathcal{H}^{\alpha}_{m}}.$
However, this means that $e\geq 2(m+1)-|\alpha|$, and then we can write
$e+c\geq 2(m+1)-|\alpha|+c=2(m+1)+1+m-\nu_{\alpha}(f),$
which in turn implies that $e+c>2(m+1)$. Thus, $e>d$, and the proof is complete. ∎
Next, we show that $Jet(f)$ determines the embedded topological type of the singularity of the Newton non-degenerate plane curve defined by $f$.
@@
Let $f,g\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. If the jet-components graph $Jet(f)=Jet(g)$, then $f$ and $g$ have the same embedded topological type.
@@
By Theorem 8, we see that $J:=Jet(f)$ determines $B^{*}:=\bigcup_{m\in\mathbf{N}_{0}}B^{*}_{m}$. In other words, $B^{*}_{m}:=B^{*}_{m}(f)=B^{*}_{m}(g)$ for all $m$. Let $R^{*}$ denote the induced subgraph of $J$ which is obtained by removing the vertices of the form $\mathcal{H}^{\gamma}_{m}$ with $\gamma\in B^{*}_{m}$ and $\nu_{\gamma}(f)>m\in\mathbf{N}_{0}$ and such that $\mathcal{H}^{\gamma}_{m}$ could not be joined with a vertex outside of $B^{*}$. Next, we define $A=\{\alpha:\exists m:\mathcal{H}^{\alpha}_{m}\in R^{*}\}$, and we note that
$\alpha\in A\Leftrightarrow\mathcal{H}^{\alpha}_{\nu_{\alpha}(f)-1}\in R^{*},\alpha\in T(f)\cap\mathbf{N}^{2}.$
Now let $T=\bigcup_{\alpha\in A}\mathbf{R}_{\geq 0}\cdot\alpha$ which is the tropical variety of both $f$ and $g$ and let $f=f_{1}...f_{r}$ and $g=g_{1}...g_{s}$ be written as product of branches. In addition, for $\alpha_{i}\in A$, let $T_{\alpha_{i}}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ denote the tropical variety of the branches $f_{1},...,f_{r_{i}},g_{1},...,g_{s_{i}}$ with $r_{i}\leq r$ and $s_{i}\leq s$. Since $f$ and $g$ are Newton non-degenerate with the same graph $J$ then there are exactly $r_{i}=s_{i}$ vertices above $\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(f)-1}=\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(g)-1}$ connected to it. Then, $r=s$ and for $1\leq i<j\leq l<\ell\leq r$ we have
$(f_{i},f_{j})_{0}=\nu_{\alpha_{i}}(f_{i})=\nu_{\alpha_{i}}(f_{j}),$
$(f_{i},f_{\ell})_{0}=\nu_{\alpha_{i}}(f_{\ell}).$
Clearly, the same holds for $g$. Now if we denote by $F=f/(f_{1}...f_{l})$ and $G=g/(g_{1}...g_{l})$, then
$l\nu_{\alpha_{i}}(f_{1})+\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)=l\nu_{\alpha_{i}}(g_{1})+\nu_{\alpha_{i}}(G),$
$\nu_{\alpha_{i}}(f_{1})=\mbox{lcm}(\alpha_{i,1},\alpha_{i,2})=\nu_{\alpha_{i}}(g_{1})$. Since $\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)$ then $\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(G)$ which holds for all $\alpha_{i}\in A$. Thus, we have shown that $f$ and $g$ have the same number of branches and the same intersection multiplicities (of their branches) as well as the same tropical varieties for branches (since $r_{i}=s_{i}$). Finally, in view of Theorem 3, $f$ and $g$ have the same embedded topological type. ∎
As we did earlier in Figures 1, 3 and 4, we can associate to $J=Jet(f)$ a special staircase walk $J_{SC}(f)$ in $\mathbf{R}^{2}_{\geq 0}$ as follows. With the same notation as in Theorem 9, and with $T_{i}=T(f_{i})$ where we assume that $T_{j}$ above $T_{i}$ for $i<j$, then we can define a road-map for this walk: every drawn $\alpha\in\mathbf{N}^{2}$ represents the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$ of $\mathcal{C}^{0}_{m}$ with $m<\nu_{\alpha}(f)$. The red arrows correspond to the irreducible components of $\mathcal{C}^{0}_{m}$ that do not vanish and continue till $\infty$. Up to this point, we distinguish the following 3 cases.
@@
    If $t=1$, then the staircase walk is identical to that shown in Figure 3 except this time, we add $r$ red arrows above $\alpha$ at each $\alpha\in T(f)$. More explicitly, if we let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$ and with the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from Theorem 1, then our walk is given by the following moves:
    • The starting point is at $(1,1)$ where $\lambda=(0,0)$,
    • move $d_{1}$ unit steps vertically followed by one horizontal step,
    • move $d_{2}$ unit steps vertically followed by one horizontal step,
    • continuing this way, we move $d_{p-1}$ unit steps vertically followed by one horizontal step,
    • our last move here consists of moving $d_{p}-1$ vertical steps to reach $(p,q)$.
    • Next, we add this time $r$ red arrow above $(p,q)$,
    • then we take $\lambda=(p,q)$ and restart again.
    Now we keep repeating the above steps for $\lambda\in L_{(p,q)}\cap\mathbf{N}_{0}^{2}$ until reaching $\lambda+(p,q)$ above which we add $r$ red arrows. It is worthy to note here that each time we move from a point $\alpha$ to either $\alpha+(0,1)$ or $\alpha+(1,0)$, we delete $\alpha$ and replace it by the new point. However, the red arrows are never deleted.
@@
    If $t=2$, we set $\sigma=T_{1}+T_{2}$; then the staircase walk obeys the following rules:
    • Start always at $(1,1)$,
    • if $\alpha\in\overset{\circ}{\sigma}$, then replace $\alpha$ with $\alpha+(0,1)$ and $\alpha+(1,0)$,
    • if $\alpha\in T_{1}$, then draw $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0)$,
    • if $\alpha$ is above $T_{2}$ then replace $\alpha$ by $\alpha+(1,0)$,
    • if $\alpha\in T_{2}$, then draw $r_{2}$ red arrows above $\alpha$ and add $\alpha+(0,1)$,
    • if $\alpha$ is below $T_{1}$ then replace $\alpha$ by $\alpha+(0,1)$.
@@
    If $t>2$, we set $T=T(f)$ and $\sigma=T_{1}+T_{t}$ then the staircase walk obeys the following rules:
    • Start always at $(1,1)$,
    • if $\T\alpha\in\sigma\backslash$ T, then replace $\alpha$ with $\alpha+(0,1)$ and $\alpha+(1,0)$,
    • if $\alpha\in T_{i}$ where $2\leq i\leq t-1$, then draw $r_{i}$ red arrows above $\alpha$ and add $\alpha+(1,0)$ with $\alpha+(0,1)$,
    • if $\alpha\in T_{1}$, then draw $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0),$
    • if $\alpha$ is above $T_{t}$ then replace $\alpha$ by $\alpha+(1,0)$,
    • if $\alpha\in T_{t}$, then draw $r_{t}$ red arrows above $\alpha$ and add $\alpha+(0,1)$,
    • if $\alpha$ is below $T_{1}$ then replace $\alpha$ by $\alpha+(0,1)$.
@@
$\mathbf{A}^{2}_{m}\cap F^{\alpha}_{ij,k}:\alpha\in T(f_{i})$ Non-vanishing $(\alpha,ij,\infty)$
Table 3. Dictionary of notation for a Newton non-degenerate $f$.
@@
For $\mathcal{C}=\{f=0\}\subset k^{2}$ a Newton plane curve singularity, we can extract the irreducible components of $\mathcal{C}^{0}_{m}$ for all $m\in\mathbf{N}$ from $J_{SC}(f)$.
@@
Consider first an edge and denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$. Now consider a non-vanishing component represented by $(\alpha,ij,\infty)$. Add on the arrow above $\alpha$, a countably infinite number of vertices. These represent the non-vanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$. The obtained graph is the jet components graph. ∎